OpenStudy (kittiwitti1):

Calculus, single-variable. Use the epsilon-delta limit definition to prove the following: (LaTeX below)

OpenStudy (kittiwitti1):

\[\huge{\lim_{x\rightarrow-2}\frac{1}{(x+2)^{2}}=\infty}\]

OpenStudy (kittiwitti1):

I'm not sure what exactly to do. What values would I put for \(\epsilon\) and \(\delta\)?

OpenStudy (kittiwitti1):

*\(\Large{\epsilon}\) and \(\Large{\delta}\) I am aware that \(\Large{\epsilon\gt0}\).

OpenStudy (welshfella):

cant remember that stuff sorry. Its been a long time...

OpenStudy (kittiwitti1):

That's okay, thanks for trying :)

OpenStudy (karim728):

do you remember your epsilon delta definition?

OpenStudy (kittiwitti1):

Yeah but how do I apply that

OpenStudy (kittiwitti1):

I can't get f(x)-L as far as I know bc L is \(\infty\)

OpenStudy (karim728):